# TAMO:Fine-Grained Root Cause Analysis via Tool-Assisted LLM Agent with Multi-Modality Observation Data

### v2

@@ III-A Data Feature Observation Tool ($\mathscr{T}_{1}$)
$q\left(x_{1},\cdots,x_{T}\mid x_{0}\right):=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right),$ (1)
@@
$q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\left(1-\alpha_{t}\right)\mathbf{I}\right),$ (2)
@@
$c_{log}=\left\{\arg\max_{e\in\mathcal{E}}\operatorname{TF-IDF}(e)\right\}\cup\left\{\arg\max_{w\in\mathcal{W}}\operatorname{TF-IDF}(w)\right\},$ (4)
@@
$\mathbf{c}_{time}=\operatorname{Embedding}\left(x_{kpi}\right)\in\mathbb{R}^{h_{{time}}},$ (5)
@@
$\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right)=\mu\epsilon_{log}\left(\mathbf{x}_{t},t,\mathbf{c}_{log}\right)+\left(1-\mu\right)\epsilon_{time}\left(\mathbf{x}_{t},t,\mathbf{c}_{time}\right),$ (7)
@@
$\mathcal{L}_{diff}=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(x_{t},t,\mathcal{C})\right\|_{2}^{2}\right],$ (8)
Through this denoising process, we obtain time series data $x^{\prime}_{ent}\in\mathbb{R}^{L\times d_{ent}}$ that integrates multimodal features for subsequent analysis, where $ent$ denotes the corresponding resource entity, $L$ denotes the length of the time window, and $d_{ent}$ denotes the dimension of the feature associated with each entity.
@@ III-B Root Cause Localization Tool ($\mathscr{T}_{2}$)
Since the interactions between microservices in the cloud native system can be described by a dependency graph, in order to fully consider the relationships between entities in the system and capture potential fault propagation pathways, we use the self-attention mechanism to construct a causal dependency graph in the frequency domain. The process can be defined as:
@@
where $W^{Q}$,$W^{K}$ are the query and key matrices respectively. Subsequently, to reduce the influence of irrelevant entities and remove redundant relationships, we introduce topological constraints into the graph generation process. These constraints can be defined as:
@@
$\xi(i,j)=LeakyRelu\left(a^{T}\left(W^{s}\cdot h^{s}_{i}\oplus W^{s}\cdot h^{s}_{j}\right)\right),$ (17)
@@
$\mathcal{L}_{BCE}=-\frac{1}{E}\sum_{i=1}^{E}\left[y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right],$ (20)
III-C Fault Classification Tool ($\mathscr{T}_{3}$)
@@
$\text{Attention}=\text{softmax}\left(\frac{zW^{\hat{Q}}\cdot z{W^{\hat{K}}}^{T}}{\sqrt{d_{k}}}\right)zW^{\hat{V}},$ (21)
where $W^{\hat{Q}}$,$W^{\hat{K}}$,$W^{\hat{V}}$ are linear projection matrices. Subsequently, we concatenate the results of the multi-head attention to obtain the latent vector representation:
$V=\text{MultiHead}\left(zW^{\hat{Q}},zW^{\hat{K}},zW^{\hat{V}}\right),$ (22)
@@
$\mathcal{L}_{wBCE}=-\frac{1}{E}\sum_{i=1}^{E}\left[\lambda_{w}\cdot y_{i}\log(\hat{y_{i}})+(1-y_{i})\log(1-\hat{y_{i}})\right],$ (24)
@@ III-D RCA Expert Agent ($\mathscr{A}$)
After the three tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ have completed tasks such as data feature alignment, root cause localization, and fault classification, the framework has achieved a relatively deep understanding of the fault. However, relying solely on these outputs makes it difficult to provide comprehensive and accurate failure analysis and remediation plans. Therefore, to further improve fault response and system recovery efficiency, we introduce the RCA expert agent $\mathscr{A}$. Using large language models, this agent synthesizes the predictions from the previous tools and incorporates the current context of cloud-native microservice architecture to offer a more comprehensive and accurate fault diagnosis and resolution strategy for site reliability engineers.
@@
  • Fault Classification Results (from $\mathscr{T}_{3}$): Providing the specific type of fault, such as network failure, hardware failure, configuration errors, etc.
  • System Context Information: Including the system microservice architecture, dependencies between services, call chains, resource usage, and other related details, which provide essential context for fault diagnosis.
In practice, we prepare the prompt in the format shown in Fig.3 as input to the model, and the expected output is a detailed fault analysis report along with corresponding solution recommendations. Such outputs assist the SRE team in quickly understanding the root cause, impact scope, and taking targeted remediation actions, thus enhancing system reliability and response efficiency.
TABLE I: The experimental results of different approaches on the tasks of root cause localization and fault classification.
Data Approach RE Localization Fault Type Classification
@@ IV-A Research Questions
  • RQ2: How effective is TAMO in Fault Classification?
@@
  • RQ4: How well does TAMO perform in real-world case studies and fault scenarios?
@@ IV-B1 Datasets
  • Dataset $A$. $A$ is a large-scale public dataset from the AIOps Challenge [^1] https://competition.AIOps-challenge.com/home/competition/1496398526429724760, which was collected through fault injection into the real-world deployed microservice system HipsterShop [^2] https://github.com/GoogleCloudPlatform/microservices-demo. The dataset primarily consists of three types of data: metrics, logs, and traces. The system uses a dynamic deployment architecture, consisting of 10 services, each with 4 pods, totaling 40 pods that are dynamically deployed across 6 nodes. The dataset includes 15 types of faults in total. Among these, 9 fault types are related to services and pods in the context of Kubernetes (K8s) containers [5], while node faults include 6 types: sudden memory pressure, disk space exhaustion, disk read/write issues, CPU pressure, and slow CPU growth.
  • Dataset $B$. $B$ is a public dataset [^3] https://doi.org/10.5281/zenodo.7615393 collected from the broadcast-style SocialNetwork system in [23], which only contains service-level entities. It includes 21 microservices and also consists of three types of data: metrics, logs, and traces. The dataset was generated by injecting faults into the system using Chaosblade, producing data that includes three types of faults: CPU resource exhaustion, network latency, and packet loss.
@@ IV-B2 Baselines
We compared TAMO with the most popular integrated multimodal methods and derivative methods from the time series domain. Specifically, we selected two state-of-the-art integrated multimodal root cause analysis methods, Eadro[23] and HolisticRCA[14], as well as four unimodal time series analysis methods (i.e., TimesNet[36], SegRNN[26], DejaVu[25], Transformer[33]). These methods were implemented using their publicly available and reproducible open-source code, and we adjusted the data format to match the required input. Furthermore, since our research task is relatively novel and includes both fault localization and classification, we incorporated the same fault classifier used in TAMO for methods that do not natively implement fault classification (such as methods [36], [23], [33], etc.) for comparison. Furthermore, we introduced LightGBM[20], a method specifically designed for fault classification tasks, as a benchmark to evaluate TAMO performance.
TABLE II: The experimental results of the ablation study conducted on our proposed TAMO framework.
Models RE Localization Fault Type Classification
@@ IV-B3 Evaluation metrics
TAMO is dedicated to locating root cause instances and identifying fault types. Based on previous research works[14],[23],[42],[15], we selected different evaluation metrics to evaluate the performance of model for these two tasks. Specifically, for the root cause localization task, we use top-k accuracy (Acc@K) as the evaluation metric, which represents the probability that the true root cause entity appears within the top-k results produced by the method. Here, we set $k=\{1,3,5\}$. Given $S$ as the set of faults, $RC$ is the ground truth root cause, and $rc[k]$ are the top-k root causes generated by the model. Therefore, Acc@K can be defined as follows:
@@
For the fault classification task, we use $Precision(Pre)$, $Recall(Rec)$, and $F1-score(F1)$ to evaluate performance. Given true positives (TP), false positives (FP), and false negatives (FN), the formulas for these metrics are as follows: $Pre=\frac{TP}{TP+FP}$,$Rec=\frac{TP}{TP+FN}$, $F1=\frac{2\cdot Pre\cdot Rec}{Pre+Rec}$. Considering that this task is a multi-label classification problem, we adopt both micro and macro metrics for a comprehensive evaluation. Micro metrics aggregate the performance across all labels, treating all instances equally, which is useful when the class distribution is imbalanced. In contrast, macro metrics compute the metric for each label independently and then average the results, giving equal weight to each label regardless of its frequency. Thus, the final performance is measured using micro-precision ($MiPr$), macro-precision ($MaPr$), micro-recall ($MiRe$), macro-recall ($MaRe$), micro-F1 score ($MiF1$), and macro-F1 score ($MaF1$).
@@ IV-B4 Implementation
We implemented TAMO using Python torch 2.4.0 and CUDA 12.1. The experiments were conducted on a Linux server equipped with an NVIDIA GeForce RTX 3090 GPU, using Python 3.8. The model was trained with the Adam optimizer [22], and all experiments used a fixed random seed. For hyperparameters, we set the batch size to 32, $\mu$ to 0.5, $\beta$ to 0.001, and the initial learning rate to 0.001. Additionally, due to the heterogeneity of system entities in Dataset $A$, which includes three levels: Pod, Service, and Node, the collected data features vary across these levels, making it challenging to train certain baselines uniformly. For such baselines, we partition Dataset $A$ into subsets composed of different entity types (e.g., $A_{p}$, $A_{s}$, and $A_{n}$) and perform separate training and testing for each subset to evaluate their performance.
@@ IV-C RQ1: Effectiveness in Root Cause Localization
Table I presents the results for root cause localization. The experimental data shows that our proposed TAMO model achieves the best or second-best performance across all datasets. Specifically, on the $A_{s}$ dataset, the score of TAMO exceeds the second-best score by at least 6%, and even in the second-best performance case, the gap with the best result is kept under 3%, clearly outperforming other competitors. Furthermore, we observe that the Eadro and Transformer models, which perform well on the B dataset, experience a significant drop in performance on heterogeneous datasets such as $A_{s}$, $A_{p}$, and $A_{n}$. This is because these two methods can only handle homogeneous entity features (e.g., in the dataset B, there is only one type of service entity). When confronted with heterogeneous entity features that need to be processed separately, their performance significantly deteriorates due to the lack of inter-entity correlation information. In contrast, both HolisticRCA and our approach model heterogeneous entities in a unified manner, maintaining strong performance even on the dataset A. Compared to HolisticRCA, we map multimodal information into time-series features and further optimize it through spatiotemporal modeling, achieving better performance in the root cause localization task.
IV-D RQ2: Effectiveness in Fault Classification
In the fault classification task, as shown in Table 2, TAMO demonstrates significant advantages in two key metrics: Micro Precision (MiPr) and Macro F1 (MaF1). Particularly on the $A_{n}$ dataset, TAMO achieves MaPr of 0.8869 and MaF1 of 0.8706, marking improvements of 24.32% and 19.85% respectively over the second-best model(HolisticRCA). TAMO stands out across all metrics such as Precision, Recall, and F1. Furthermore, it exhibits remarkable performance in the classification of complex fault types at the node level. In addition, TAMO excels in the precision metric, with MiPr reaching 0.7164 and 0.7182 on $A_{s}$ and $A_{p}$, respectively, surpassing the best baseline model by 8. 09% and 12. 69%, significantly reducing the risk of misclassification. Although MiPr on the smaller dataset $B$ is lower than the baseline model LightGBM, TAMO still outperforms LightGBM in terms of MaPr and F1 metrics. This suggests that TAMO performs more evenly across different classes, achieving better classification results for minority classes with fewer samples.

### v4

@@ III-A Multi-modality Alignment Tool ($\mathscr{T}_{1}$)
$q\left(x_{1},\cdots,x_{T}\mid x_{0}\right):=\prod_{t=1}^{T}q\left(x_{t}\mid x_{t-1}\right).$ (1)
@@
$q\left(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}\right)=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1},\left(1-\alpha_{t}\right)\mathbf{I}\right).$ (2)
@@
$c_{log}=\left\{\arg\max_{e\in\mathcal{E}}\operatorname{TF-IDF}(e)\right\}\cup\left\{\arg\max_{w\in\mathcal{W}}\operatorname{TF-IDF}(w)\right\}.$ (4)
@@
$c_{time}=\operatorname{Embedding}\left(x_{kpi}\right)\in\mathbb{R}^{h_{{time}}}.$ (5)
@@
$\epsilon_{\theta}\left(\mathbf{x}_{t},t,\mathcal{C}\right)=\mu\epsilon_{log}\left(\mathbf{x}_{t},t,\mathbf{c}_{log}\right)+\left(1-\mu\right)\epsilon_{time}\left(\mathbf{x}_{t},t,c_{time}\right).$ (7)
@@
$\mathcal{L}_{diff}=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(x_{t},t,\mathcal{C})\right\|_{2}^{2}\right].$ (8)
Through this denoising process, we obtain time series data $x^{\prime}_{ent}\in\mathbb{R}^{L\times d_{ent}}$ that integrates multimodal features for subsequent analysis, where entent denotes the corresponding resource entity, $L$ denotes the length of the time window, and $d_{ent}$ denotes the dimension of the feature associated with each entity. This generative alignment avoids the semantic misalignment inherent in feature concatenation, yielding a more coherent and temporally consistent multimodal representation for downstream analysis ($\mathscr{T}_{2}$ and $\mathscr{T}_{3}$).
@@ III-B Root Cause Localization Tool ($\mathscr{T}_{2}$)
Since the interactions between microservices in the cloud-native systems can be described by a dependency graph, in order to fully consider the relationships between entities in the system and capture potential fault propagation pathways, we use the self-attention mechanism to construct a causal dependency graph in the frequency domain. The process can be defined as:
@@
where $W^{Q}$$W^{K}$ are the query and key matrices respectively. Subsequently, to reduce the influence of irrelevant entities and remove redundant relationships, we introduce topological constraints into the graph generation process. These constraints can be defined as:
@@
$\xi(i,j)=LeakyRelu\left(a^{\top}\left(W^{s}\cdot h^{s}_{i}\oplus W^{s}\cdot h^{s}_{j}\right)\right),$ (17)
@@
[…]
III-C Fault Types Classification Tool ($\mathscr{T}_{3}$)
@@
$\text{Attention}=\text{softmax}\left(\frac{zW^{\hat{Q}}\cdot z{(W^{\hat{K}}})^{\top}}{\sqrt{d_{k}}}\right)zW^{\hat{V}}$, (21)
where $W^{\hat{Q}}$$W^{\hat{K}}$$W^{\hat{V}}$ are linear projection matrices. Subsequently, we concatenate the results of the multi-head attention to obtain the latent vector representation:
[…]
@@
[…]
@@ III-D RCA Expert Agent ($\mathscr{A}$)
After the three tools $\mathscr{T}_{1}$-$\mathscr{T}_{3}$ have completed tasks such as multi-modality alignment, root cause localization, and fault types classification, the framework has achieved a relatively deep understanding of the fault. However, relying solely on these outputs makes it difficult to provide comprehensive and accurate failure analysis and remediation plans. Therefore, to further improve fault response and system recovery efficiency, we introduce the RCA expert agent $\mathscr{A}$. Using large language models, this agent synthesizes the predictions from the previous tools and incorporates the current context of cloud-native microservice architecture to offer a more comprehensive and accurate fault diagnosis and resolution strategy for site reliability engineers.
@@
  • Fault Types Classification Results (from $\mathscr{T}_{3}$): Providing the specific type of fault, such as network failure, hardware failure, configuration errors, etc.
  • System Context Information: Includes the microservice architecture (e.g., service roles, dependencies, and call chains), deployment topology, and resource usage patterns. This context, often derived from system documentation, guides the agent in understanding operational relationships and diagnosing faults.
In practice, we prepare the prompt in the format shown in Fig.3 as input to the model, and the expected output is a detailed fault analysis report along with corresponding solution recommendations. Such outputs assist the engineers in quickly understanding the root cause, impact scope, and taking reasonable remediation actions, thus enhancing system reliability and response efficiency.
TABLE II: Experimental results of different approaches on root cause localization and fault classification. Best results are bolded, and second-best results are underlined
Data Approach Root Cause Localization Fault Types Classification
@@ IV-A Research Questions
  • RQ2: How effective is TAMO in Fault Types Classification?
@@
  • RQ4: What are the training and inference efficiency of TAMO, and how do hyperparameters affect its performance?
  • RQ5: How well does TAMO perform in real-world case studies and fault scenarios?
@@ IV-B1 Datasets
  • Dataset $A$. AA is a large-scale public dataset from the AIOps Challenge [^1] https://competition.AIOps-challenge.com/home/competition/1496398526429724760, which is collected through fault injection into the real-world deployed microservice system HipsterShop [^2] https://github.com/GoogleCloudPlatform/microservices-demo. The dataset primarily consists of three types of data: metrics, logs, and traces. The system uses a dynamic deployment architecture, consisting of 10 services, each with 4 pods, totaling 40 pods that are dynamically deployed across 6 nodes. The dataset includes 15 types of faults in total. Among these, 9 fault types are related to services and pods in the context of Kubernetes (K8s) containers [4], while node faults include 6 types: sudden memory pressure, disk space exhaustion, disk read/write issues, CPU pressure, and slow CPU growth.
  • Dataset $B$. $B$ is a public dataset [^3] https://doi.org/10.5281/zenodo.7615393 collected from the broadcast-style SocialNetwork system in [23], which only contains service-level entities. It includes 21 microservices and also consists of three types of data: metrics, logs, and traces. The dataset is generated by injecting faults into the system using Chaosblade, producing data that includes three types of faults: CPU resource exhaustion, network latency, and packet loss.
@@ IV-B2 Baselines
We compare TAMO with the most popular integrated multimodal methods and derivative methods from the time series domain. Specifically, we select two state-of-the-art integrated multimodal root cause analysis methods, Eadro[23] and HolisticRCA[14], as well as four unimodal time series analysis methods (i.e., TimesNet[42], SegRNN[28], DejaVu[27], Transformer[39]). These methods are implemented using their publicly available and reproducible open-source code, and we adjust the data format to match the required input. Furthermore, since our research task is relatively novel and includes both fault localization and classification, we incorporate the same fault classifier used in TAMO for methods that do not natively implement fault classification (such as methods [23], [39], [42], etc.) for comparison. Furthermore, we introduce LightGBM[20], a method specifically designed for fault classification tasks, as a benchmark to evaluate TAMO performance.
TABLE III: The experimental results of the ablation study conducted on our proposed TAMO framework.
Models Root Cause Localization Fault Types Classification
@@ IV-B3 Evaluation metrics
TAMO is dedicated to locating root cause instances and identifying fault types. Based on previous research works[14][15][23][48], we select different evaluation metrics to evaluate the performance of model for these two tasks. Specifically, for the root cause localization task, we use top-k accuracy (Acc@K) as the evaluation metric, which represents the probability that the true root cause entity appears within the top-k results produced by the method. Here, we set $k=\{1,3,5\}$. Given $S$ as the set of faults, $RC$ is the ground truth root cause, and $rc[k]$ are the top-k root causes generated by the model. Therefore, Acc@K can be defined as follows:
@@
For the fault classification task, we use $Precision(Pre)$, $Recall(Rec)$, and F1score(F1)F1-score(F1) to evaluate performance. Given true positives (TP), false positives (FP), and false negatives (FN), the formulas for these metrics are as follows: $Pre=\frac{TP}{TP+FP}$$Rec=\frac{TP}{TP+FN}$, $F1=\frac{2\cdot Pre\cdot Rec}{Pre+Rec}$. Considering that this task is a multi-label classification problem, we adopt both micro and macro metrics for a comprehensive evaluation. Micro metrics aggregate the performance across all labels, treating all instances equally, which is useful when the class distribution is imbalanced. In contrast, macro metrics compute the metric for each label independently and then average the results, giving equal weight to each label regardless of its frequency. Thus, the final performance is measured using micro-precision ($MiPr$), macro-precision ($MaPr$), micro-recall ($MiRe$), macro-recall ($MaRe$), micro-F1 score ($MiF1$), and macro-F1 score ($MaF1$).
@@ IV-B4 Implementation
We implement TAMO using Python torch 2.4.0 and CUDA 12.1. The experiments are conducted on a Linux server equipped with an NVIDIA GeForce RTX 3090 GPU, using Python 3.8. The model is trained with the Adam optimizer [22], and all experiments use a fixed random seed. For hyperparameters, we set the batch size to 32, $\mu$ to 0.5, $\beta$ to 0.001, epochepoch to 200, and the initial learning rate to 0.001. Considering the requirements of TAMO for system contextual awareness and temporal consistency, data cleaning is performed on datasets A and B before training. This process remove noisy and misaligned multimodal records and realigned log and trace data to ensure temporal consistency and enhance the quality of input observation data for $\mathscr{T}_{1}$. Additionally, considering the pre-training needs of $\mathscr{T}_{1}$, an extra dataset of observation data from normal runtime operations is constructed. This dataset is used for $\mathscr{T}_{1}$ pre-training to enable learning of normal system patterns.
For training, we adopt a sequential and modular strategy to ensure stable learning across the tool ensemble. We first pre-train tool $\mathscr{T}_{1}$ using normal operational data from Datasets A and B, leveraging the diffusion framework to learn robust representations of multi-modal system observations. Once $\mathscr{T}_{1}$ is pre-trained, its parameters are frozen and it serves as a fixed feature extractor for subsequent stages. Tool $\mathscr{T}_{2}$ is then trained using the embeddings generated by $\mathscr{T}_{1}$ to perform root cause localization. After $\mathscr{T}_{2}$ is trained, both $\mathscr{T}_{1}$ and $\mathscr{T}_{2}$ are kept frozen while training $\mathscr{T}_{1}$-generated embeddings and the causal graphs from $\mathscr{T}_{2}$ as input for fault classification. Finally, all components are jointly fine-tuned for 5 epochs with a small learning rate of 0.0001 to refine the overall pipeline. This staged training ensures effective knowledge transfer and prevents instability during end-to-end optimization.
Additionally, due to the heterogeneity of system entities in Dataset $A$, which includes three levels: Pod, Service, and Node. The collected data features vary across these levels, making it challenging to train certain baselines uniformly. For such baselines, we partition Dataset $A$ into subsets composed of different entity types (e.g., $A_{p}$, $A_{s}$, and $A_{n}$) and perform separate training and testing for each subset to evaluate their performance. Finally, the symbol ’-’ is used in the results table II to indicate cases where a method failed to achieve meaningful classification performance (e.g., accuracy close to 0) on different datasets.
Figure 4: Hyperparameter sensitivity analysis across batch size, $\mu$, regularization strength ($\beta$), and hidden dimension.
@@ IV-C RQ1: Effectiveness in Root Cause Localization
Table II presents the results for root cause localization. The experimental data shows that our proposed TAMO model achieves the best or second-best performance across all datasets. Specifically, on the $A_{s}$ dataset, the score of TAMO exceeds the second-best score by at least 6%, and even in the second-best performance case, the gap with the best result is kept under 3%, clearly outperforming other baselines. Furthermore, we observe that the Eadro and Transformer models, which perform well on the B dataset, experience a significant drop in performance on heterogeneous datasets such as $A_{s}$, $A_{p}$, and $A_{n}$. This is because these two methods can only handle homogeneous entity features (e.g., in the dataset B, there is only one type of service entity). When confronted with heterogeneous entity features that need to be processed separately, their performance significantly deteriorates due to the lack of inter-entity correlation information. In contrast, both HolisticRCA and our approach model heterogeneous entities in a unified manner, maintaining strong performance even on the dataset A. Compared to HolisticRCA, we map multimodal information into time-series features and further optimize it through spatiotemporal modeling, achieving better performance in the root cause localization task.
IV-D RQ2: Effectiveness in Fault Types Classification
In the fault classification task, as shown in Table II, TAMO demonstrates significant improvments in two key metrics: Micro Precision (MiPr) and Micro F1 (MiF1). Particularly on the $A_{n}$ dataset, TAMO achieves a MiPr of $0.8718$ and MiF1 of $0.8831$, outperforming the second-best model (HolisticRCA) by $24.99\%$ and $19.85\%$, respectively, and ranks first across all evaluation metrics. With MiPr reaching $0.7164$ and $0.7182$ on $A_{s}$ and $A_{p}$, respectively, TAMO outperforms the best baseline by $8.09\%$ and $12.69\%$, significantly reducing the misclassification. Furthermore, we observe that transformer-based models achieve relatively strong performance on relatively simpler dataset B. However, their performance degrades significantly on the large-scale, complex system dataset A, indicating that metric-only transformer architectures struggle to capture the intricate, multi-modal dependencies inherent in large-scale distributed systems.
Figure 5: Comparision of training time and inference time of TAMO with Eadro.